what electrostatic force would exist between a point charge with a charge of 5.35x10 -7C and a second point charge with a charge of -7.73x10 -6C if the two charges were separated by a distance of 4.11x10 -2?


1
Expert's answer
2021-02-25T12:37:30-0500

Here,

"q_1=5.35\times10^{-7}C"

"q_2=-7.73\times10^{-6}C"

"d=4.11\times 10^{-2}m"


For electrostatic force, "F_E= k\dfrac{q_1 q_2}{d^2}"

and "k = 9\times 10^9 Nm^2/C^2"

So, "F_E=(9\times 10^9\dfrac{Nm^2}{C^2})\dfrac{(5.35\times10^{-7}C)(-7.73\times10^{-6}C)}{(4.11\times10^{-2}m)^2}"

"=\dfrac{-372.1995}{16.8921}=-22.03N"


So, magnitude of electrostatic force "|F_E|=" 22.03N

and negative sign shows that there is attraction between the two point charges.